OpenStudy (anonymous):

The length of a rectangle is 2 more than 3 times the width. If the perimeter is 100, find the length and width of the rectangle. If l = length and w = width, which of the following systems could be used to solve the problem? l = 3w + 2 and 2l + 2w = 100 l = 2w + 3 and 2l + 2w = 100 l + 2 = 3w and 2l + 2w = 100

OpenStudy (anonymous):

L=2+3W 2L + 2W = 100 the only thing you have to think about is the equation of the perimeter

OpenStudy (anonymous):

since it's a rectangle, it's 2W +2L = P... in this case P was 100